OpenStudy (18jonea):

How many moles are in 2.75 x 10^24 molecules of SO3 @sweetburger

OpenStudy (sweetburger):

the conversion factor for moles to molecules is 1 mole = 6.022 x 10^23 molecules therefore you complete this process

OpenStudy (18jonea):

ok how would i do that?

OpenStudy (18jonea):

?

OpenStudy (18jonea):

i think it is 1.16

OpenStudy (18jonea):

sorry i meant 4.57

OpenStudy (18jonea):

@sweetburger

OpenStudy (18jonea):

is that right?

OpenStudy (18jonea):

@sweetburger

OpenStudy (sweetburger):

yes ur answer was correct sorry i had to step away from teh keyboard

OpenStudy (unklerhaukus):

\[n =\frac N{N_A} =\frac{2.75\times10^{24}\,[\text {molecule}\cdot\text S]}{6.02\times10^{23}\,[\text {molecule}/\text{mol}]} =4.57\,[\text{mol}\cdot\text S]\]